# Beyond Homophily: Community Search on Heterophilic Graphs

### v1

@@ V-B Global weight renormalization
$\mathbf{w}^{(k)}=\sigma\!\left(U\Delta\Sigma^{(k)}V^{\!\top}\mathbf{1}\right),\quad X^{(k)}=\mathbf{w}^{(k)}\odot X^{(k)},$ (20)
@@ VI Methodology Phase ii: Online Searching
In the online search phase, we introduce two methods for efficiently retrieving communities from precomputed embeddings. The first method leverages a Signed Community Search (SCS) algorithm that operates on a positive signed graph constructed from node embeddings, and incorporates a teleportation step to effectively handle sparse connections by allowing direct jumps to nodes highly similar to the query. The second method, Adaptive Community Score (ACS), selects candidate nodes based on their embedding similarity to the query and then ranks them using a homophily-adaptive scoring function. This function combines similarity and direct-connection rewards or penalties, automatically adjusting to the estimated homophily level of the graph. Both methods are designed to handle the challenges presented by heterophilic graphs, ensuring accurate and meaningful community retrieval.
@@ VI-A Online Search via SCS
We first solve the absence of edge signs in heterophilic graphs by using the learned node embeddings from the AdaptCS encoder to infer the sign of each edge. For each edge $(u,v)$ in the original graph, we compute the cosine similarity between their normalized embeddings; if the similarity exceeds a threshold $\tau$, we treat $(u,v)$ as a positive edge. This yields a pruned, semantically meaningful subgraph that avoids introducing spurious edges between unconnected nodes. Community search then proceeds via a BFS exploration over the positive graph, starting from the query node. Due to the predominant number of negative edges and sparsity in heterophilic graphs, BFS may frequently reach a dead-end before the desired community size $\mathcal{K}$ is reached. To address this, we use a teleportation mechanism: whenever BFS is exhausted but $|\mathcal{C}_{q}|<\mathcal{K}$, we select the unvisited node most similar to the query node (according to the learned embedding similarity) and resume BFS from there. This guarantees that the community is always extended with nodes most relevant to the query, maintaining semantic coherence.
@@ VI-B Online Search via ACS
Input :  Query node $q$, embeddings $H\in\mathbb{R}^{n\times h}$, adjacency matrix $A$, community size $\mathcal{K}$, similarity-weighting parameter $\tau$, hyperparameters: $\lambda_{\text{bonus}}$, $\lambda_{\text{penalty}}$, $\alpha$
Output : Community $\mathcal{C}_{q}\subseteq V$
1
2Estimate global homophily $h_{edge}$ from random sample
3 $w_{\text{bon}}\leftarrow(1-\tau)h_{edge}\cdot\lambda_{\text{bonus}}$
4 $w_{\text{pen}}\leftarrow(1-\tau)(-(1-h_{edge})\cdot\lambda_{\text{penalty}})$
5 Compute cosine similarities $S_{qu}$ for all $u$
6 Select candidate set $C$ as top-$(\alpha\cdot\mathcal{K})$ most similar nodes to $q$
7 for each $u$ $\in$ $C$ do
8 if $u$ is adjacent to $q$ then
9    $\text{score}[u]=S_{qu}+(w_{\text{bon}}$ if $h_{edge}\geq 0.5$ else $w_{\text{pen}})$
10   end if
11 else
12    $\text{score}[u]=S_{qu}$
13   end if
14
15 end for
16Sort candidates in descending order of score
Return $\mathcal{C}_{q}=$ top-$\mathcal{K}$ nodes from sorted list (plus $q$)
Algorithm 1 Community Search via Adaptive Community Score (ACS)
ACS search procedure. Algorithm 1 first estimates the global homophily ratio and selects a candidate set by semantic similarity to the query node. For each candidate, it computes a homophily-adaptive score that combines similarity and, if directly connected to the query, a topology-aware bonus or penalty. The $\mathcal{K}$ highest-scoring nodes (plus the query itself) form the returned community. This method is robust across varying homophily levels and avoids the pitfalls of methods that rely solely on topology or similarity.
@@ VII Theoretical Analysis
VII-A Distance Awareness Guarantees HDD
Prior work ACM [ACM_2022] proved that, when the number of communities satisfies $c=2$, the diversification distinguishability (DD) metric achieves the optimal value $\mathrm{DD}=1$ under ACM filtering. Building upon this result, we show that our distance-aware distinctive-hop filtering guarantees high-order diversification distinguishability $\mathrm{HDD}=1$, given when the number of communities $c>2$.
@@ VII-A Distance Awareness Guarantees HDD
Building on the assumption $X=Z\in\mathbb{R}^{n\times c}$ in ACM, we set the attention bank as $B=Z$, meaning each class corresponds to a one-hot filter. Recall the per-class attention:
$\alpha_{\mathrm{class}}=\mathrm{softmax}(ZB)=\mathrm{softmax}(ZZ^{\top}),$
where $ZZ^{\top}$ is a $N\times N$ binary matrix: $(ZZ^{\top})_{uv}=1$ iff $u$ and $v$ have the same label, 0 otherwise. Therefore, for each node $v$, $\alpha_{\mathrm{class},v}$ is a one-hot vector on $c_{v}$ (the class of $v$).
The adaptive feature weight for node $v$ is
$w_{\mathrm{class},v}=\alpha_{\mathrm{class},v}B^{\top}=\text{one-hot}(c_{v})\in\mathbb{R}^{c},$
which means, for every node $v$, $w_{\mathrm{class},v}$ is $1$ only at the $c_{v}$-th dimension, and 0 elsewhere. Now consider the $k$-hop distinctive-hop feature: $H_{v}^{(k)}=\hat{L}^{(k)}Z,$ where $\hat{L}^{(k)}$ is the exact-$k$-hop Laplacian matrix. After applying per-class bank weighting, we have: $\widetilde{H}_{v}^{(k)}=H_{v}^{(k)}\odot w_{\mathrm{class},v}.$
Since $w_{\mathrm{class},v}$ is one-hot at $c_{v}$, this step sets all non-class features to zero for node $v$. That is, even if $v$ aggregates information from heterophilic neighbors, only the entry corresponding to $v$’s class remains after weighting. The final embedding is then: $H_{v}=\sum_{k=1}^{K}\alpha_{\mathrm{hop},v}^{(k)}\widetilde{H}_{v}^{(k)}.$ Consider two nodes $u,v$: If $c_{u}=c_{v}=c$, then both $H_{u}$ and $H_{v}$ are nonzero only at index $c$. Their similarity: $\langle H_{u},H_{v}\rangle\geq 0,$ since they only have support at the same class coordinate, and any positive aggregation increases similarity. If $c_{u}\neq c_{v}$, then $H_{u}$ and $H_{v}$ are nonzero at disjoint class coordinates. Thus,
$\langle H_{u},H_{v}\rangle=0.$
This holds regardless of the presence of heterophilic noisy neighbors, as the per-class bank attention removes non-class contribution for each node. Therefore, for any $v$,
$\mathrm{Mean}_{u}\{\,\langle H_{v},H_{u}\rangle:c_{u}=c_{v}\,\}$ $\geq 0,$
$\mathrm{Mean}_{u}\{\,\langle H_{v},H_{u}\rangle:c_{u}\neq c_{v}\,\}$ $=0.$
This satisfies the high-order diversification distinguishability (HDD) definition, so $\mathrm{HDD}_{H_{HP}}(G)=1$ for any $c>2$.
VII-B Triangle-support lower bound for adaptively retained edges
Proof.
Because $\hat{A}$ is symmetric and $(u,v)\in E$, $\begin{aligned} (\hat{A}^{2})_{uv}=\hat{A}_{uu}\hat{A}_{uv}+\hat{A}_{uv}\hat{A}_{vv}+\sum_{w\in CN(u,v)}\hat{A}_{uw}\hat{A}_{wv}.\end{aligned}$ Using $\hat{A}_{uv}=1/\sqrt{\hat{d}_{u}\hat{d}_{v}}$, $\hat{A}_{uu}=1/\hat{d}_{u}$, $\hat{A}_{vv}=1/\hat{d}_{v}$, and $\hat{A}_{uw}\hat{A}_{wv}=\frac{1}{\sqrt{\hat{d}_{u}\hat{d}_{v}}}\cdot\frac{1}{\hat{d}_{w}}$ for $w\in CN(u,v)$, we obtain
$(\hat{A}^{2}-\hat{A})_{uv}=\frac{1}{\sqrt{\hat{d}_{u}\hat{d}_{v}}}\left(\frac{1}{\hat{d}_{u}}+\frac{1}{\hat{d}_{v}}+\sum_{w\in CN(u,v)}\frac{1}{\hat{d}_{w}}-1\right).$
The prefactor is positive, so the retention condition $(\hat{A}^{2}-\hat{A})_{uv}>0$ is equivalent to
$\sum_{w\in CN(u,v)}\frac{1}{\hat{d}_{w}}\;>\;1-\frac{1}{\hat{d}_{u}}-\frac{1}{\hat{d}_{v}}\;=\;T(u,v).$
Next, for any common neighbor $w\in CN(u,v)$, because we work with $A+I$, the node $w$ is incident to at least the edges $(w,u)$, $(w,v)$, and the self-loop $(w,w)$, hence $\hat{d}_{w}\geq 3$, which implies $\frac{1}{\hat{d}_{w}}\;\leq\;\frac{1}{3}\ \text{for all }w\in CN(u,v).$ Summing over $CN(u,v)$ gives
$\sum_{w\in CN(u,v)}\frac{1}{\hat{d}_{w}}\;\leq\;\frac{|CN(u,v)|}{3}\;=\;\frac{\mathrm{supp}(u,v)}{3}.$
Combining above equations yields
$\frac{\mathrm{supp}(u,v)}{3}\;>\;T(u,v)\quad\Rightarrow\quad\mathrm{supp}(u,v)\;>\;3\,T(u,v).$
Since $\mathrm{supp}(u,v)$ is an integer, we conclude
$\mathrm{supp}(u,v)\;\geq\;\big\lfloor 3\,T(u,v)\big\rfloor+1,$
which match the claimed bound. ∎
TABLE I: Dataset statistics
Datasets #nodes #edges #features #classes $h_{edge}$
Cora 2,708 5,429 1,433 7 0.8100
CiteSeer 3,327 4,732 3,703 6 0.7362
Photo 7,650 119,081 745 8 0.8272
Computers 13,752 245,861 767 10 0.7772
DBLP 17,716 52,867 1,639 4 0.8279
CS 18,333 81,894 6,805 15 0.8081
PubMed 19,717 44,338 500 3 0.8024
Reddit 232,956 116M 602 41 0.7817
Cornell 183 295 1,703 5 0.5669
Texas 183 309 1,703 5 0.4106
Wisconsin 251 499 1,703 5 0.4480
Chameleon 2,277 31,396 2,325 5 0.2299
Squirrel 5,201 198,423 2,089 5 0.2221
Film 7,600 33,544 931 5 0.3750
Roman 22,662 16,463 300 18 0.0469
Flickr 89,250 449,878 500 7 0.3195
TABLE II: Effectiveness evaluation of different datasets
@@ VII-B Triangle-support lower bound for adaptively retained edges
VII-C Time Complexity Analysis
Let $n=|V|$, $m=|E|$, $d$ be the input feature dimension, $h$ the hidden size, $k$ the number of hops, $r$ the SVD rank, $t$ the number of training epochs, $Q$ the number of queries, and $\mathcal{K}$ the target community size.
Offline pre-computation. We compute an $r$-rank SVD of the $n\times n$ sparse adjacency (with $m$ nonzeros) in $O(r\times m\times I)$ time (where $I$ is the number of Lanczos iterations). Projecting features requires $O(n\times r\times d)$. For each of the $k$ distinctive hops, forming the diagonal correction and multiplying by $U$ and the projected features takes $O(n\times r\times d)$ per hop, yielding $O\bigl(r\times m\times I+(k+1)\times n\times r\times d\bigr).$
Offline training. Each epoch performs $k$ hops of two linear transforms of cost $O(n\times h^{2})$ each, plus an optional sparse multiplication of cost $O(m\times h)$ for structure-aware attention. Accounting for forward and backward passes, one epoch costs $O\bigl(k\times n\times h^{2}+m\times h\bigr),$ and $t$ epochs give $O\bigl(t\times(k\times n\times h^{2}+m\times h)\bigr)$.
Signed community score (SCS). This builds the positive graph costs $O(m\times d)$, and each query runs a BFS with teleport in $O(n+m)$. Over $Q$ queries this is $O\bigl(m\times d+Q\times(n+m)\bigr).$
Adaptive community score (ACS). It first samples $s=\min(m,m_{s})$ edges in $O(s\times d)$, then for each query computes a full cosine similarity vector in $O(n\times d)$, selects the top($\alpha.\mathcal{K}$) candidates in $O(n\log n)$, and scores $\mathcal{K}$ nodes with neighbor checks in $O(\mathcal{K}\log d_{\max})$. Thus, per query it costs $O\bigl(n\times d+n\log n+\mathcal{K}\log d_{\max}\bigr),$ and overall $O\bigl(s\times d+Q\times(n\times d+n\log n+\mathcal{K}\log d_{\max})\bigr)$.
@@ VIII-A Experimental Setup
Datasets. We evaluate AdaptCS on 16 real-world graphs with varying levels of edge homophily. Eight datasets exhibit high homophily, including Cora, CiteSeer, PubMed [cora_yang2016revisiting], Amazon-Computers, Amazon-Photo, Coauthor-CS [shchur2018pitfalls], DBLP [tang2008arnetminer], and Reddit [reddit]. The others show low homophily (heterophilic) patterns, including Cornell, Wisconsin, Texas, Film [geom-gcn], Chameleon (corrected), Squirrel (corrected), Roman [platonovcritical], and Flickr [factorLP_2020_SIGMOD]. Statistics are summarized in Table I.
@@
Baseline models. We compare AdaptCS against three categories of baselines: (i) Algorithm-based methods including $k$-core [k-core], CTC [k-truss2015], and $k$-clique [k-clique]; (ii) H-based Label propagation (LP) methods, evaluated in both vanilla and attributed settings, serving as non-parametric embedding-free baselines; and (iii) ML-based community search (ML-CS) models, including ICSGNN [ics-gnn], QDGNN [qdgnn], COCLEP [coclep], and ComDF [chen2025communitydf]. Since these ML-CS models were originally developed under the homophily assumption, we extend each with three heterophily-oriented variants: (1) Tanh – following FAGCN [FAGCN21], we replace the unbounded ReLU with Tanh to allow negative aggregation; (2) ACM – we integrate Adaptive Channel Mixing [ACM_2022], which separates low- and high-frequency signals; and (3) ALT – we apply Adaptive Local Tuning [ALT_KDD_2023], a dynamic weighting mechanism for cross-class propagation. Each ML baseline is therefore evaluated under the three mentioned extensions to ensure fair and comprehensive comparison across homophilic and heterophilic graphs.
@@
Evaluation metrics. The evaluation of identified communities is conducted through F1-score [ics-gnn, qdgnn]. The F1-score balances precision and recall, offering a measure of how well the identified community matches the ground truth. The true data is established as the target community label, with the labels of the identified nodes serving as the predicted data. To evaluate the efficiency, the model training time and online querying time are recorded across different models. All results are averaged across 50 randomly selected queries to ensure the quality of the evaluation process.
@@
Implementation details. In the experiment, We run AdaptCS for 100 epochs with early stopping. AdaptCS uses a 5-hop receptive field and 512 hidden units by default. For large datasets like Reddit, we use a 3-hop receptive field and 128 hidden units to accommodate their larger scale. The learning rate is set as 0.01. The community size $\mathcal{K}$ is dataset-dependent and may vary according to user needs. Specifically, we set $\mathcal{K}=30$ for small datasets, $150$ for medium-sized datasets, and $1000$ for large datasets. The experiments are run on a machine with Intel Xeon 6248R CPU, Nvidia A5000 GPU, and 512GB memory. The code is available at Github  [^1] https://github.com/SimonQS/AdaptCS.

### v2

@@ V-B Global weight renormalization
$\mathbf{w}^{(k)}=\sigma\!\left(U_{r}\Delta\Sigma_{r}^{(k)}V_{r}^{\!\top}\mathbf{1}\right),\quad X^{(k)}=\mathbf{w}^{(k)}\odot X^{(k)},$ (22)
@@ VI Methodology Phase ii: Online Searching
In the online search phase, we introduce two methods for efficiently retrieving communities from precomputed embeddings. The first method leverages a Signed Community Search (SCS) algorithm that operates on a positive signed graph constructed from node embeddings, and incorporates a teleportation step to effectively handle sparse connections by allowing direct jumps to nodes highly similar to the query. Adaptive Community Score (ACS) selects candidate nodes based on their embedding similarity to the query and then ranks them using a homophily-adaptive scoring function. Both methods are designed to handle the challenges presented by heterophilic graphs, ensuring accurate and meaningful community retrieval.
@@ VI-A Online Search via SCS
We first solve the absence of edge signs in heterophilic graphs by using the learned node embeddings from the AdaptCS encoder to infer the sign of each edge. For each edge $(u,v)$ in the original graph, we compute the cosine similarity between their normalized embeddings; if the similarity exceeds a threshold $\tau$, we treat $(u,v)$ as a positive edge. This yields a pruned, semantically meaningful subgraph that avoids introducing spurious edges between unconnected nodes. Community search then proceeds via a BFS exploration over the positive graph, starting from the query node. Due to the predominant number of negative edges and sparsity in heterophilic graphs, BFS may frequently reach a dead-end before the desired community size $\mathcal{K}$ is reached. To address this, we use a teleportation mechanism: whenever BFS is exhausted but $|\mathcal{C}_{q}|<\mathcal{K}$, where $|\mathcal{C}_{q}|$ is the size of the current community, we select the unvisited node most similar to the query node (according to the learned embedding similarity) and resume BFS from there. Algorithm pseudocode is disclosed in Appendix A online [39].
@@ VI-B Online Search via ACS
ACS search procedure. ACS first estimates the global homophily ratio and selects a candidate set by semantic similarity to the query node. For each candidate, it computes a homophily-adaptive score that combines similarity and, if directly connected to the query, a topology-aware bonus or penalty. The $\mathcal{K}$ highest-scoring nodes (plus the query itself) form the returned community. This method is robust across varying homophily levels and avoids the pitfalls of methods that rely solely on topology or similarity. Algorithm pseudocode is disclosed in Appendix A online [39].
@@ VII Theoretical Analysis
VII-A Distance Awareness Suppresses Similarity Flip
Theorem 2 (Distance-aware hop separation suppresses similarity flip).
Recall the HND gap $\Phi_{u}(\cdot)$ in Definition IV.1. Let $z_{u}\in\{1,\dots,c\}$ be the class label of node $u$, and let $\{p_{j}\}_{j=1}^{c}\subset\mathbb{R}^{d}$ be a prototype bank with $\|p_{j}\|\leq P$ for all $j$. Given a root representation $r_{u}\in\mathbb{R}^{d}$ and exact-hop representations $h_{u}^{(1)},h_{u}^{(2)}\in\mathbb{R}^{d}$, our fusion first forms class weights $\alpha^{class}_{u,j}$, then computes hop-attention weights $\alpha^{hop}_{u,k}$. Assume $\|h_{u}^{(1)}\|,\|h_{u}^{(2)}\|\leq B$ for all $u$. If there exists an informative hop (here $k=2$) such that $\Phi_{u}(h^{(2)})\geq\delta$ for all $u$, and the above bank-conditioned hop attention selects this hop for each node $u$ in the large-temperature regime, then for sufficiently large $(\beta_{c},\beta_{h})$ we have $\Phi_{u}(h)>0$ for all $u$, hence $\mathrm{HND}(G)=1$. Moreover, a sufficient condition for hop-$2$ selection is the existence of margins $\gamma,\eta>0$ such that $\langle r_{u},p_{z_{u}}\rangle\geq\max_{j\neq z_{u}}\langle r_{u},p_{j}\rangle+\gamma$ and $\langle h_{u}^{(2)},p_{z_{u}}\rangle\geq\langle h_{u}^{(1)},p_{z_{u}}\rangle+\eta$ for all $u$.
Remark (Effect of feature noise). When the input features are noisy, e.g., $x_{u}=x_{u}^{\star}+\xi_{u}$ with $\|\xi_{u}\|\leq\epsilon$. Such bounded noise perturbs inner products by at most $O(\epsilon)$, and therefore reduces the margins (e.g., $\gamma,\eta$) and the informative-hop gap (e.g., $\delta$) by a bounded amount. As long as the perturbed margins and gap remain positive, the conclusion of Theorem 2 still holds.
@@ VII-A Distance Awareness Suppresses Similarity Flip
We verify hop-$2$ selection under the stated sufficient margin conditions and then conclude HND.
Step 1 (root margin $\Rightarrow w_{u}\approx p_{z_{u}}$). Let $z_{u}$ be the true class of $u$ and recall $\alpha^{class}_{u,j}\propto\exp(\beta_{c}\langle r_{u},p_{j}\rangle)$. If $\langle r_{u},p_{z_{u}}\rangle\geq\max_{j\neq z_{u}}\langle r_{u},p_{j}\rangle+\gamma$, then for any $j\neq z_{u}$,
$\frac{\alpha^{class}_{u,j}}{\alpha^{class}_{u,z_{u}}}=\exp\!\big(\beta_{c}(\langle r_{u},p_{j}\rangle-\langle r_{u},p_{z_{u}}\rangle)\big)\leq e^{-\beta_{c}\gamma}.$
Thus $\sum_{j\neq z_{u}}\alpha^{class}_{u,j}\leq(c-1)e^{-\beta_{c}\gamma}$ and
$\|w_{u}-p_{z_{u}}\|\leq 2P\sum_{j\neq z_{u}}\alpha^{class}_{u,j}\leq 2P(c-1)e^{-\beta_{c}\gamma}.$
so $w_{u}=p_{z_{u}}+O\!\big(P(c-1)e^{-\beta_{c}\gamma}\big)$.
Step 2 (hop margin $\Rightarrow\alpha^{hop}_{u,2}\approx 1$). For $k\in\{1,2\}$, since $\|h_{u}^{(k)}\|\leq B$ and $\|w_{u}-p_{z_{u}}\|\leq 2P(c-1)e^{-\beta_{c}\gamma}$,
$\big|\langle h_{u}^{(k)},w_{u}\rangle-\langle h_{u}^{(k)},p_{z_{u}}\rangle\big|\leq 2BP(c-1)e^{-\beta_{c}\gamma}.$
If additionally $\langle h_{u}^{(2)},p_{z_{u}}\rangle\geq\langle h_{u}^{(1)},p_{z_{u}}\rangle+\eta$, then
$\langle h_{u}^{(2)},w_{u}\rangle-\langle h_{u}^{(1)},w_{u}\rangle\geq\eta-4BP(c-1)e^{-\beta_{c}\gamma}.$
For sufficiently large $\beta_{c}$ such that $4BP(c-1)e^{-\beta_{c}\gamma}\leq\eta/2$, we have $\langle h_{u}^{(2)},w_{u}\rangle\geq\langle h_{u}^{(1)},w_{u}\rangle+\eta/2$. Since $\alpha^{hop}_{u,k}\propto\exp(\beta_{h}\langle h_{u}^{(k)},w_{u}\rangle)$, this implies
$\Delta_{u}:=\langle h_{u}^{(2)},w_{u}\rangle-\langle h_{u}^{(1)},w_{u}\rangle\geq\eta/2$
$\Rightarrow\alpha^{hop}_{u,2}=\frac{1}{1+e^{-\beta_{h}\Delta_{u}}}\geq\frac{1}{1+e^{-\beta_{h}\eta/2}}.$
Therefore, using $h_{u}=\sum_{k\in\{1,2\}}\alpha^{hop}_{u,k}h_{u}^{(k)}$ and $\|h_{u}^{(k)}\|\leq B$,
$\|h_{u}-h_{u}^{(2)}\|\leq 2B(1-\alpha^{hop}_{u,2})\leq 2Be^{-\beta_{h}\eta/2}.$
Step 3 (stability of the HND gap). For inner-product similarity (or $\epsilon$-stabilized cosine), the similarity $S(u,v)$ and hence the induced gap $\Phi_{u}(\cdot)$ vary continuously on the bounded set $\{\|h\|\leq B\}$. Thus the perturbation $\|h_{u}-h_{u}^{(2)}\|\leq 2Be^{-\beta_{h}\eta/2}$ implies
$\Phi_{u}(h)\geq\Phi_{u}(h^{(2)})-O\!\big(e^{-\beta_{h}\eta/2}\big).$
Using the informative-hop condition $\Phi_{u}(h^{(2)})\geq\delta$, we obtain
$\Phi_{u}(h)\geq\delta-O\!\big(e^{-\beta_{h}\eta/2}\big).$
Having $\beta_{h}$ sufficiently makes the error smaller than $\delta$, so $\Phi_{u}(h)>0$ for all $u$, which satisfy the HND condition. ∎
VII-B Time Complexity Analysis
Query efficiency comparison. SCS costs $O(n+m)$ per query, while ACS costs $O(nh+M\log M)$, dominated by the similarity scan. Full analysis is in Appendix A online [39].
ICSGNN/QDGNN/COCLEP are query-driven: each query requires a query-conditioned forward propagation over the graph. Notably, the original ICSGNN implementation is fully online (i.e., it trains for each query without a separate offline stage); for fair comparison, we restructured it into an offline-training plus online-search pipeline similar to QDGNN. However, in these methods, each query still requires $k$-layer inference, giving a shared per-query complexity $O\!\big(k(mh+nh^{2})\big).$
CommunityDF iteratively refine on a query-induced subgraph $G_{s}=(V_{s},E_{s})$: it first computes a distribution (e.g., $O(|E_{s}|+|V_{s}|\log|V_{s}|)$ via PPR or $O(L(|E_{s}|h+|V_{s}|h^{2}))$ via GCN scoring) and then runs $T_{c}$ diffusion steps. Thus, its per-query cost scales with $T_{c}$ and can be summarized as $O\!\big(|E_{s}|+|V_{s}|\log|V_{s}|\big)\;+\;O\!\big(T_{c}\,L(|E_{s}|h+|V_{s}|h^{2})\big),$ where $T_{c}\ll T$ is the reduced diffusion steps in their optimization.
Overall, AdaptCS/ACS is query-decoupled: it avoids per-query graph-wide $k$-layer inference and only performs a similarity scan plus candidate ranking. In contrast, query-driven methods incur $O(k(mh+nh^{2}))$ per query due to graph propagation, and CommunityDF further adds iterative refinement whose cost grows with $T_{c}$ on the query-induced subgraph.
TABLE I: Effectiveness evaluation of different datasets
@@ VIII-A Experimental Setup
Datasets. We evaluate AdaptCS on 16 real-world graphs with varying levels of homophily. Eight datasets exhibit high homophily, including Cora, CiteSeer, PubMed [42], Amazon-Computers, Amazon-Photo, Coauthor-CS [43], DBLP [44], and Reddit [45]. The others show heterophilic patterns, including Cornell, Wisconsin, Texas, Film [35], Chameleon (corrected), Squirrel (corrected), Roman [46], and Flickr [26]. Statistics are summarized in Appendix B online [39].
@@
Baseline models. We compare AdaptCS against three categories of baselines: (i) Algorithm-based methods including $k$-core [7], CTC [47], and $k$-clique [11]; (ii) H-based Label propagation (LP) methods, evaluated in both vanilla and attributed settings, serving as non-parametric embedding-free baselines; and (iii) ML-based community search (ML-CS) models, including ICSGNN [13], QDGNN [4], COCLEP [14], and ComDF [34]. Since these ML-CS models were developed under the homophily assumption, we evaluate each baseline under a consistent set of heterophily-oriented extensions to ensure fair comparison across both homophilic and heterophilic graphs, including Tanh [27], MixHop [36], Geom-GCN [35], ACM [28], and ALT [29]. We adopt the original hyperparameter settings reported in the respective baseline papers.
@@
Evaluation metrics. The evaluation of identified communities is conducted through F1-score [13, 4]. The F1-score balances precision and recall, offering a measure of how well the identified community matches the ground truth. The true data is established as the target community label, with the labels of the identified nodes serving as the predicted data. To evaluate the efficiency, the model training time and online querying time are recorded across different models. All results are averaged across 50 randomly selected queries.
@@
Implementation details. The ground-truth community of a query is defined as the set of nodes sharing the same label as the query node. Following the standard supervised settings [33, 4], we split labeled nodes into train/val/test as 60%/20%/20%. The training set label is only used to compute the loss and update model parameters. During validation, the parameters are frozen, and the test set remains untouched to prevent information leakage. All query nodes are randomly sampled from the test set only. We run AdaptCS for 100 epochs with early stopping. AdaptCS uses a 5-hop receptive field and 512 hidden units by default. We use a 3-hop receptive field and 128 hidden units for large datasets to accommodate their scale. The attention-bank fusion is used for hop channel mixing. The learning rate is set as 0.01. The community size $\mathcal{K}$ is dataset-dependent and may vary according to user needs. We set $\mathcal{K}=30$ for small datasets, $150$ for medium-sized datasets, and $1000$ for large datasets. The experiments are run on a machine with Intel Xeon 6248R CPU, Nvidia A5000 GPU, and 512GB memory. The code is available at Github  [^1] https://github.com/SimonQS/AdaptCS.
VIII-B Effectiveness of AdaptCS
Table I benchmarks our approach on 16 graphs, where AdaptCS-I is the vanilla model that performs distance- and frequency-aware aggregation on the full adjacency; AdaptCS-II adds the proposed low-rank SVD compression, which keeps hop-distinct features in the latent space and thus avoids memory overflow. For the online search phase, we compare two method: the Signed Community Search (SCS) and Adaptive Community Score (ACS). “-” indicates that the method exceeds the GPU memory limit or fails to finish within the 12-hour wall-clock time. The rightmost column, Ave +/-, reports the average F1-score difference of each method relative to the best AdaptCS-II (ACS) result on average across all datasets.
Observation 3 – Homophily-oriented GNN backbones remain limited. Among the four ML-CS backbones, QDGNN achieves the best overall average (with its strongest variant under Geom-GCN), followed by ICSGNN and ComDF, while COCLEP consistently underperforms. With the expanded set of five heterophily-oriented extensions/backbones (Tanh, MixHop, Geom-GCN, ACM, ALT), we observe that Geom-GCN and Tanh tend to yield the most consistent improvements for message-passing baselines, while ACM remains a competitive and stable alternative; ALT provides mixed gains and can even degrade performance for some backbones (with ALT also incurring higher overhead). Nevertheless, even the best-configured baseline still falls short of AdaptCS.